OpenStudy (anonymous):

Simplify the expression. 12p - 10a + 4 - 4p -6

OpenStudy (anonymous):

NEED the answer!

OpenStudy (anonymous):

12p-4p = 8p 4-6 = -2 So: 8p=10a-2 p= (5/4)a-(1/5)

OpenStudy (anonymous):

But what is the answer to this problem?

OpenStudy (anonymous):

The general idea of simplifying, when you have no other information, is to eliminate all common factors (eg. 12p-4p = 8p), and juggle to isolate the expression for one variable. The equation you posted originally, was: 12p - 10a + 4 - 4p -6 Which is really the same as: 12p - 10a + 4 - 4p -6=0 12p-4p-10a+4-6=0 8p-10a-2=0 8p=10a+2 (divide by 8 to isolate for p:) p=(1/8)(10a+2) => p= (5/4)a + (1/4) (typo in my last reply)

OpenStudy (anonymous):

I know but what is the answer.

OpenStudy (anonymous):

The answer? You haven't provided enough information for me to help you any further, I'm afraid. Sure the equation is correct? If that's all the information you've got, that is the answer.

OpenStudy (anonymous):

This is what it says..... Simplify the expression. 11p - 7q - 12 + 9q - 5r + 8

OpenStudy (anonymous):

Oops wrong one/

OpenStudy (anonymous):

Well, it says "simplify". That's what I've done. Isolating for one variable is as simple as it gets. Alternatively you can write the expression as 8p-10a-2